A HOPF INDEX THEOREM FOR FOLIATIONS

VICTOR BELFI, EFTON PARK, AND KEN RICHARDSON

Abstract. We formulate and prove an analog of the Hopf Index Theorem for Riemannian foliations. We compute the basic Euler characteristic of a closed Riemannian manifold as a sum of indices of a non-degenerate basic vector field at critical leaf closures. The primary tool used to establish this result is an adaptation to foliations of the Witten deformation method.